Supply Club, Inc., sells a variety of paper products, office supplies, and other products used by businesses and individual consumers. During July 2021 it started a loyalty program through which qualifying customers can accumulate points and redeem those points for discounts on future purchases. Redemption of a loyalty point reduces the price of one dollar of future purchases by 20% (equal to 20 cents). Customers do not earn additional loyalty points for purchases on which loyalty points are redeemed. Based on past experience, Supply Club estimates a 80% probability that any point issued will be redeemed for the discount. During July 2021, the company records $211,500 of revenue and awards 146,875 loyalty points. The aggregate stand-alone selling price of the purchased products is $211,500. Eighty percent of sales were cash sales, and the remainder were credit sales.

Required: 1. & 2. Prepare Supply Club’s journal entry to record July and August sales. During August, customers redeem loyalty points on $94,000 of merchandise. Seventy-five percent of those sales were for cash, and the remainder were credit sales. (Do not round intermediate calculations. If no entry is required for a transaction/event, select "No journal entry required" in the first account field.)

pleaase I need help with this question
Answer #1

No.

Date

Account titles and explanation

Debit

Credit

1

July 31, 2021

Cash (211500*80%)

169200

Accounts receivable (211500*20%)

42300

Revenue

21150

Deferred revenue

190350

2

August 31, 2021

Cash (94000*(1-20%)*75%)

56400

Accounts receivable (94000*(1-20%)*25%)

18800

Deferred revenue

16920

Revenue

92120

Loyalty Points = 146875*20%*80% = 23500

Allocation:

Loyalty Points: 23500 / (23500+211500) = 0.1 * 211500 = 21150

Product Sales: 211500 / (23500+211500) = 0.9 * 211500 = 190350

Deferred Revenue:

146875 points issued * 80% expected to be redeemed = 117500 points to be redeemed

94,000 points redeemed / 117500 expected = 80% redeemed

$21150 deferred revenue * 80% = $16920 Revenue earned